# SVGFusion: Scalable Text-to-SVG Generation via Vector Space Diffusion

### v1

@@ 3.2 Vector-Pixel Fusion VAE
Sampling vector graphics in one-dimensional discrete vector spaces poses significant challenges for language models, as it requires handling complex dependencies across sparse and high-dimensional textual representations, making it difficult to capture the intricate structure and continuity inherent in vector graphics. Our goal is to learn the neural representation of vector graphics by mapping vector geometric features to the latent space. Unlike previous methods [21, 3, 58, 61], our approach overcomes two primary limitations: 1) the difficulty of capturing visual information in vector graphics through implicit latent representations, which often leads to suboptimal quality in decoded SVG shapes; and 2) the challenges associated with training and sampling from an unstructured and under-constrained latent space. To achieve this, as shown in Fig. 2(b), we propose a Vector-Pixel Fusion Variational Autoencoder (VP-VAE) that learns the representations from both pixel features and vector primitives. Given the requirement to model serialized vector primitives, we employ Transformer-based architectures for both our encoders and decoders, leveraging recent findings that highlight the superiority of this approach [55, 3, 31].
@@
Vector-Pixel Fusion Encoding. Our proposed VP-VAE encoder, $\mathcal{E}(\cdot)$, accepts two types of inputs: pixel embeddings extracted from SVG renderings and SVG embeddings derived from SVG codes. This design aligns visual features with geometric embedding features during the coding phase. The encoder architecture of the VP-VAE comprises one cross-attention layer followed by $L$ self-attention layers. Specifically, the pixel embeddings initially query the SVG embeddings through the cross-attention layer. The resulting tokens then pass through the subsequent self-attention layers. The final output latent code $\bm{z}$ from the encoder are designed to encapsulate both visual and geometric features.
SVG Decoding. The decoder of VP-VAE employs a similar structure to the encoder, but without the cross-attention layer. The latent code $\bm{z}$ is decoded into the SVG embeddings by the decoder. Finally, the reconstructed SVG embeddings are fed into a linear layer to predict the command and parameter logits.
VP-VAE Objective. We first quantify the differences between the predicted primitive $\mathrm{E}_{i}$ and the ground-truth primitive $\hat{\mathrm{E}}_{i}$ via cross-entropy loss. In addition to reconstruction losses associated with both components, the latent space undergoes regularization via a KL divergence loss. The final loss of VP-VAE is presented in Eq. 1:
$\mathcal{L}_{\mathrm{VAE}}=\mathcal{L}_{\mathrm{CE}}(\mathrm{E}_{i},\hat{\mathrm{E}}_{i})+\lambda_{\mathrm{KL}}\mathcal{L}_{\mathrm{KL}}$ (1)
where $\mathcal{L}_{\mathrm{KL}}$ encourages the latent code $\bm{z}$ to follow a Gaussian distribution $\mathcal{N}(\bm{0},\bm{I})$, and $\lambda_{\mathrm{KL}}$ denotes the weight of KL regularization.
@@ 3.3 Vector Space Diffusion Transformer
Figure 3: The architecture of our VP-DiT Block. A cross-attention module is integrated into each block to inject textual conditions.
SVGFusion adopts the DiT [31] as the base architecture and perform the diffusion process in the vector latent space. To allow textual features $f(\mathcal{P})$ to interact with vector features, similar to [4], we incorporate a multi-head cross-attention layer into the VS-DiT block. It is positioned between the self-attention layer and the feed-forward layer so that the model can flexibly interact with the text embedding extracted from the language model. Our architecture supports flexible stacking of $N_{b}$ VS-DiT blocks, enabling models with varied parameter scales that allow SVGFusion to adaptively expand with increasing data volume, thus enhancing scalability with data growth. Given an input SVG through VP-VAE: $\bm{z}=\mathcal{E}(\mathcal{G})$, diffused inputs $\bm{z}_{t}=\alpha_{t}\bm{z}+\sigma_{t}\epsilon,\epsilon\sim\mathcal{N}(\bm{0},\bm{I})$ are constructed; $\alpha_{t}$ and $\sigma_{t}$ define a noise schedule, parameterized via a diffusion-time $t$. Following LDM [38], our VP-DiT model predicts the noise $\epsilon$ of the noisy latent representation $\bm{z}_{t}$ at time $t$, conditioned on text prompt $y$. The objective is shown in Eq. 2:
$\mathcal{L}_{\mathrm{ldm}}=\mathbb{E}_{\mathcal{E}(\bm{x}),y,t,\epsilon\sim\mathcal{N}(\bm{0},\bm{I})}\left[\left\|\epsilon_{\phi}(\bm{z}_{t},t,y)-\epsilon\right\|_{2}^{2}\right]$ (2)
where $t\sim\mathcal{U}(0,1)$. When training the diffusion model, we randomly zero the $y$ with a probability of 10% to use classifier-free guidance [13], which can improve the quality of conditional generation during inference.
@@ 4 Experiments
Overview. This section covers the key aspects of our model implementation, followed by a description of the dataset collection and preprocessing pipeline. We then evaluate the performance of SVGFusion in generating high-quality SVGs by comparing it to state-of-the-art methods, both quantitatively and qualitatively in Sec. 4.2 and Sec. 4.3. The evaluation also includes an architectural analysis and ablation studies to highlight the contributions of individual model components. Implementation Details. To ensure consistency across all SVG data, we adopted relative positional coordinates. The model parameters are initialized randomly and optimized using the AdamW optimizer (with $\beta_{1}=0.9$, $\beta_{2}=0.95$) at an initial learning rate of $3\times 10^{-4}$. The learning rate is warmed up over the first 2000 steps and then decayed to $1.5\times 10^{-5}$ following a cosine schedule. Additionally, we applied weight decay of 0.1 for regularization and constrain gradients by clipping their norms to a maximum value of 2.0. Besides, we normalized the input SVG embeddings into the $[-1,1]$ range to stabilize the training process. We utilized Transformers as the fundamental building block for VP-VAE. Both the encoders and decoders are based on the Transformer architecture, consisting of 4 layers, a feed-forward dimension of 512, and an embedding size of $d_{E}=256$. To investigate scaling trends, we trained VP-DiT models at three different sizes: 0.16B, 0.37B, and 0.76B parameters. We trained VP-VAE for 100k steps using a total batch size of 64 across 2 A800 GPUs, requiring approximately one day. Subsequently, leveraging the trained VP-VAE, we trained VP-DiT for 500k steps with a batch size of 128 on 8 A800 GPUs, taking approximately three days. In the sampling phase, we used DDIM [43] to de-noise 100 steps; dpm-solver [22] also supports 20-step de-noising for more efficient sampling.
@@
Table 1: Quantitative Comparison of SVGFusion vs. State-of-the-Art Text-to-SVG Methods. Our method uses dpm-solver [22] for 20-step denoising.
4.1 SVG Data Collection and Representation
Figure 4: Samples from our collected SVG dataset. The collected data includes a variety of SVG element types.
We collected a large-scale SVG dataset for model training and evaluation, which includes 240k high-quality SVGs with emoji and icon style. Further details of this new dataset are provided in Supplementary Sec. D. SVG data collected from the Internet is often noisy, and using it directly for learning can hinder the model’s representation capabilities. About half of the data in an SVG file is redundant for visual rendering, including: 1) temporary data from vector editing applications, 2) non-optimal structural representations, and 3) unused or invisible graphic elements. To address this, we propose a preprocessing pipeline that reduces SVG file size losslessly, targeting four key aspects: elements, attributes, paths, and outputs. A detailed example is shown in Fig. S6 of Supplementary. The collected dataset is highly diverse, as depicted in Fig. 4. In terms of visual complexity, it includes primitives that vary from simple to intricate. Regarding primitive representation, the dataset includes both Bézier curves expressed using <path> and basic shapes represented by <circle> and <rect>, etc. In terms of color representation, the dataset contains both black-and-white graphics and vibrant, harmonious color images. Additionally, in terms of semantics, it covers a wide range of subjects, including people, animals, objects, and symbols.
Figure 5: Qualitative Comparison of SVGFusion and Existing Text-to-SVG Methods. The target SVGs are in the emoji style. We use prompt modifiers for the optimization-based approach to encourage the appropriate style: ”minimal flat 2D vector icon, emoji icon, lineal color, on a white background, trending on ArtStation.” Note that although the visual quality of results generated by optimization-based methods is high, these methods face challenges in decomposing the SVGs for further editing. We encourage readers to zoom into this figure for a better illustration.
@@ 4.2 Quantitative Evaluation
Comparison results are presented in Table 1. The methods are categorized into two groups: optimization-based methods (top section of Table 1) and language model-based methods (middle section of Table 1). It is evident that our SVGFusion method surpasses other text-to-SVG methods across all evaluation metrics. This demonstrates the superiority of SVGFusion in generating vector graphics that are more closely aligned with text prompts and human preferences. Notably, compared to optimization-based methods, SVGFusion significantly reduces the time cost, enhancing its practicality and user-friendliness.
@@ 4.3 Qualitative Evaluation
Figure 5 presents a qualitative comparison between SVGFusion and existing text-to-SVG methods. The results are aligned with the quantitative results discussed in the previous section. Specifically, the methods Evolution [52], CLIPDraw [8], DiffSketcher [63], VectorFusion [18], LIVE [24]+VectorFusion [18], SVGDreamer [64], and Word-as-Img [17] are optimization-based, using a differentiable renderer [19] to backpropagate gradients to vector parameters. Evolution [52] and CLIPDraw [8] utilize CLIP [34] as the image prior, while DiffSketcher [63], VectorFusion [18], SVGDreamer [64], and Word-as-Img [17] adopt T2I diffusion as the image prior. Despite their visual advantages, differentiable renderer-based methods often produce intertwined vector primitives, diminishing SVG editability. Iconshop [61], Claude 3.5 [29] and GPT4o [1] rely on language models, generate decoupled vector primitives but produce overly simplistic content.
@@
Figure 6: Diffusion Process. The diffusion process is visualized during model inference.
@@
In contrast, our approach not only ensures a clear SVG hierarchy but also maintains a balance between visual appeal and complexity. Figure 6 shows the process of denoising a vector graphic over timesteps. Starting with initial random, chaotic strokes, the elements of the SVG, such as vector coordinates, are progressively denoised at each timestep. After a period of denoising, the correct content is gradually restored while maintaining the visual harmony of the generated SVG.
As illustrated in Fig. 1(a) and Fig. 7, our SVGFusion can generate SVGs using only the necessary primitives, such as <circle> and <rect>. It allows for flexibility in the design process: one can either start by sketching the general shape and then add local details, or begin with a specific part and gradually add elements to complete the SVG. This approach coincides with the design process typically used by human designers when creating SVGs.
Comparison with Large Language Model. As illustrated in Fig. 5, we also compare our proposed SVGFusion with existing state-of-the-art approaches that directly generate SVGs using Large Language Models (LLMs). Claude 3.5 and GPT-4 are currently recognized as two of the best LLMs, but their performance in SVG generation is not particularly outstanding. In most cases, they can only use simple shapes to roughly assemble objects, but the positioning of each element lacks harmony. As a result, the overall shapes are overly simplistic, and the visual effects are less satisfactory. Regarding color design, these LLMs struggle to apply colors accurately to each part of the SVG, leading to color schemes that are often neither harmonious nor reasonable. In terms of semantic expression, the SVG code generated by LLMs is too simple to fully capture the meaning conveyed by the input text description. In contrast, our proposed SVGFusion method produces more balanced and harmonious results in terms of shape selection, color matching, and semantic representation. In Supplementary Sec. B, we provide more comparison of our method with language model-based methods.
@@
Editbility of SVGFusion Results. Figure 8 shows the editability of the SVG generated by our SVGFusion. Since the SVGs we generate have a clean and concise structure, we can easily edit the properties of the primitives, such as their color attributes. For instance, the rocket we generated can be changed from blue to red, or even transformed into a cyber-metal style, simply by adjusting the color attributes. Furthermore, the capability of edibility empowers users to efficiently reuse synthesized vector elements and create new vector compositions. As illustrated in the second example of Fig. 8, our method composes a new vector graphic by replacing the torch with a umbrella.

### v2

@@ 3.2 Vector-Pixel Fusion VAE
The Vector-Pixel Fusion VAE (VP-VAE) is designed to learn a latent space for vector graphics with a transformer-based VAE architecture, which includes an encoder and a decoder. To effectively capture both structural and visual features, we introduce Vector-Pixel Fusion Encoding. This method integrates information from the SVG code and its rendered image. Additionally, to better emulate human design logic, we design a sequence modeling strategy. This strategy enables the model to learn how vector graphics are incrementally constructed, ensuring the resultant SVGs with more reasonable construction.
Vector-Pixel Fusion Encoding. As shown in Fig. 4, to effectively capture both the geometric structure and visual appearance of vector graphics, the VP-VAE encoder $\mathcal{E}(\cdot)$ takes two types of embeddings as the input: (1) SVG embeddings $\bm{E}_{\mathrm{emb}}\in\mathbb{R}^{N\times D_{e}}$. (2) Pixel embeddings $\bm{E}_{\mathrm{pix}}\in\mathbb{R}^{N\times D_{p}}$. The pixel embeddings are obtained by using a pretrained DINOv2 [31] to extract high-level visual features from the rendered images. To align these embeddings in the latent space, we first project them onto a common dimension $D_{z}$ using separate linear layers: $\bm{E}_{\mathrm{emb}}^{\prime}=\bm{W}_{\mathrm{proj}}^{e}\bm{E}_{\mathrm{emb}}^{\top}$, $\bm{E}_{\mathrm{pix}}^{\prime}=\bm{W}_{\mathrm{proj}}^{p}\bm{E}_{\mathrm{pix}}^{\top}$, where $\bm{W}_{\mathrm{proj}}^{e}\in\mathbb{R}^{D_{z}\times D_{e}},\bm{W}_{\mathrm{proj}}^{p}\in\mathbb{R}^{D_{z}\times D_{p}}$. The transformed embeddings are then subjected to a cross-attention layer, where the SVG embeddings act as queries ($Q$), while the pixel embeddings act as keys ($K$) and values ($V$). This enables the model to effectively integrate geometric and visual features, enhancing its representation of complex vector structures. Finally, the fused representation is processed through the subsequent self-attention layers and mapped to a latent distribution, from which the latent variable $\bm{z}$ is sampled.
Rendering Sequence Modeling. To enhance the construction of the generated SVGs, we introduce a Rendering Sequence Modeling strategy that enables the model to learn the creation logic of human designers. This strategy involves demonstrating to the model how an SVG is constructed from scratch by observing changes in both SVG codes and their corresponding rasterizations. During training, we represent the SVG as a progressive sequence of drawing steps. Each step incrementally adds new primitives to the SVG, simulating the progressive construction along the batch dimension. As illustrated with the cat example in Fig. 2, each SVG is processed into a batch of data of size $B$, where each sample corresponds to a different stage of creation, ranging from initial primitives to the complete SVG. This allows the model to observe the process of SVG creation within a single training iteration, thereby fostering a deeper understanding of the logical layers involved in SVG.
@@
The Architecture of VP-VAE. The encoder of the VP-VAE consists of one cross-attention layer followed by $L$ self-attention layers. Specifically, the SVG embeddings initially serve as queries to the pixel embeddings through the cross-attention layer. The tokens resulting from this interaction are then processed by the subsequent self-attention layers. The final output, latent code $\bm{z}$, from the encoder is designed to encapsulate both visual and geometric features of the SVG.
The decoder of the VP-VAE mirrors the structure of the encoder but omits the cross-attention layer. Given a latent code $\bm{z}$, the decoder reconstructs its corresponding SVG embeddings: $\bm{E}_{\mathrm{rec}}=\mathcal{D}(\bm{z})$, where $\mathcal{D}(\cdot)$ represents the decoder network. To obtain the final SVG representation, we apply two separate decoding processes: 1) Coordinate Mapping: The reconstructed embeddings containing normalized coordinates $C_{\text{rec}}$ (ranging from $[-1,1]$) are mapped back to the canvas coordinate system: $C_{\text{coord}}=\nicefrac{{(C_{\text{rec}}+1)}}{{2}}\times\mathrm{V}$, where $\mathrm{V}$ indicates the predefined canvas size. 2) Element & Command Mapping: The SVG elements and commands are recovered from their respective embeddings using a learned embedding-to-token mapping. Specifically, the reconstructed embeddings are passed through an embedding layer, which predicts the discrete SVG elements and commands, effectively reconstructing the full SVG structure. These steps ensure that the decoded SVG faithfully preserves both the structural and geometric information encoded in the latent space.
Figure 4: Illustration of the Vector-Pixel Fusion Encoding. The VP-VAE encoder integrates the SVG embeddings ($Q$) with pixel embeddings ($K$, $V$) using a cross-attention layer. After processing through $L$ self-attention layers, the encoded features are mapped to a latent space, where the mean and standard deviation are computed for a probabilistic representation. A latent variable $\bm{z}$ is sampled using the reparameterization trick and then passed to the decoder for further processing. For clarity, the batch dimension $B$ has been omitted.
VP-VAE Objective. To ensure accurate reconstruction of vector primitives, we measure the discrepancy between the predicted primitive $\mathrm{E}_{i}$ and the ground-truth primitive $\hat{\mathrm{E}}_{i}$ using the mean squared error (MSE) loss. This loss encourages the model to generate primitives that closely match the original input. Beyond reconstruction, we impose regularization on the latent space using Kullback-Leibler (KL) divergence, which constrains the learned latent distribution to approximate a standard Gaussian prior. This prevents overfitting and ensures smooth, continuous latent representations, which are crucial for generating diverse and coherent vector structures. The final VP-VAE loss is formulated as:
$\mathcal{L}_{\mathrm{VAE}}=\mathcal{L}_{\mathrm{MSE}}(\mathrm{E}_{i},\hat{\mathrm{E}}_{i})+\lambda_{\mathrm{KL}}\mathcal{L}_{\mathrm{KL}}$ (1)
where $\mathcal{L}_{\mathrm{KL}}$ measures the divergence between the learned latent distribution and the prior Gaussian distribution $\mathcal{N}(\bm{0},\bm{I})$. The KL term encourages the latent space to remain compact and structured, facilitating smooth interpolation between vector primitives.
@@ 3.3 Vector Space Diffusion Transformer
The VP-VAE effectively learns a latent space tailored for vector graphics representation. Building on this foundation, SVGFusion leverages DiT [32] as its core architecture and performs the diffusion process directly in the vector latent space. To facilitate interaction between textual features and vector representations, we introduce a multi-head cross-attention layer into the VS-DiT block, inspired by [5]. A detailed description of the VS-DiT architecture is provided in Supplementary Sec. C. During training, for a given input SVG, we first derive its latent representation using VP-VAE: $\bm{z}=\mathcal{E}(\mathcal{G})$. The diffusion process then takes place within this latent space, where the noisy latent variable $\bm{z}_{t}$ is generated as follows: $\bm{z}_{t}=\alpha_{t}\bm{z}+\sigma_{t}\epsilon,\epsilon\sim\mathcal{N}(\bm{0},\bm{I})$ where $\alpha_{t}$ and $\sigma_{t}$ define the noise schedule, parameterized by the diffusion time $t$. Following Latent Diffusion Models (LDM) [38], our VS-DiT model predicts the noise $\epsilon$ in the noisy latent representation $\bm{z}_{t}$, conditioned on a text prompt $y$. The training objective is formulated as:
$\mathcal{L}_{\mathrm{ldm}}=\mathbb{E}_{\mathcal{E}(\bm{z}),y,t,\epsilon\sim\mathcal{N}(\bm{0},\bm{I})}\left[\left\|\epsilon_{\phi}(\bm{z}_{t},t,y)-\epsilon\right\|_{2}^{2}\right]$ (2)
where $t\sim\mathcal{U}(0,1)$. We randomly set $y$ to zero with a probability of 10% to apply classifier-free guidance [15] during training, which enhances the quality of conditional generation during inference.
@@ 4 Experiments
Figure 5: Qualitative Comparison of SVGFusion and Existing Text-to-SVG Methods. The target SVGs are in the emoji style. We use prompt modifiers for the optimization-based approach to encourage the appropriate style: “minimal flat 2D vector icon, emoji icon, lineal color, on a white background, trending on ArtStation.” Note that although the visual quality of results generated by optimization-based methods is high, these methods face challenges in decomposing the SVGs for further editing.
4.1 SVGX-Dataset
We introduce SVGX-Dataset, a large-scale SVG dataset designed for model training and evaluation. It comprises 240,000 high-quality emoji and icon-style SVGs sourced from Twemoji-Color-Font [54], Noto-Emoji [10], FluentUI-Emoji [27], SVG-Repo [62], and Reshot [61].
The dataset spans a diverse range of visual complexities, incorporating both Bézier curves (expressed using <path>) and basic geometric shapes (<circle>, <rect>, etc.). It includes high-quality, human-designed SVGs in black-and-white and color, covering various semantic categories such as people, animals, objects, and symbols. Examples are shown in Fig. S1 in the Supplementary.
SVGs collected from the Internet often contain noise, which can hinder model performance. We identify three common issues: (1) temporary data generated by vector editing software, (2) suboptimal structural representations, and (3) unused or invisible graphical elements.
Although these elements do not affect the visual appearance, they introduce inefficiencies in training. To address this, we implement a cleaning pipeline that removes redundant elements, refines coordinate precision, and resizes canvases to $128\times 128$. This process significantly reduces file sizes while preserving visual fidelity. An example of the cleaning pipeline is illustrated in Fig. S3 in the Supplementary.
As shown in Fig. S2, we analyze frequently occurring words in dataset names and descriptions, visualize word distributions via a word cloud, and compare data before and after cleaning. This analysis highlights the dataset’s well-structured nature, supporting effective model training. Further details on dataset collection, preprocessing, and analysis are provided in Supplementary Sec. B.
@@ 4.1 SVGX-Dataset
Table 1: Quantitative Comparison of SVGFusion vs. State-of-the-Art Text-to-SVG Methods. Top: Optimization-based methods; Middle: language-model-based mehtods; Bottom: Three model variants of our proposed SVGFusion. Our method uses dpm-solver [24] for 20-step denoising.
@@ 4.2 Quantitative Evaluation
Comparison results are presented in Table 1. The methods are categorized into two groups: optimization-based methods (top section of Table 1) and language-model-based methods (middle section of Table 1). It is evident that our SVGFusion method surpasses other text-to-SVG methods across all evaluation metrics. This demonstrates the superiority of SVGFusion in generating vector graphics that are more closely aligned with text prompts and human preferences. Notably, compared to optimization-based methods, SVGFusion significantly reduces the time cost, enhancing its practicality and user-friendliness.
@@ 4.3 Qualitative Evaluation
Figure 5 presents a qualitative comparison between SVGFusion and existing text-to-SVG methods. The results are aligned with the quantitative results discussed in the previous section. Specifically, the optimization-based methods, including Evolution [52], CLIPDraw [9], DiffSketcher [65], VectorFusion [20], LIVE [26]+VectorFusion [20], SVGDreamer [68], and Word-as-Img [19], use a differentiable renderer [21] to backpropagate gradients to vector parameters. Evolution [52] and CLIPDraw [9] utilize CLIP [35] as the image prior, while DiffSketcher [65], VectorFusion [20], SVGDreamer [68], and Word-as-Img [19] adopt T2I diffusion as the image prior. Despite their visual advantages, optimization-based methods often produce intertwined vector primitives, diminishing SVG editability. Language-model-based methods, such as Iconshop [63], Claude3.5-sonnet [2], GPT4o-latest [1] and DeepSeek-R1 [11], rely on language models, and can generate decoupled vector primitives but overly simplistic content.
@@
As illustrated in Fig. 1(a) and Fig. 6, our SVGFusion can generate SVGs using only the necessary primitives, such as <circle> and <rect>, ensuring a compact and structured representation. It allows for flexibility in the design process: one can either start by sketching the general shape and then add local details, or begin with a specific part and gradually add elements to complete the SVG. This coincides with the process of creating SVGs by human designers.
@@
Comparison with Large Language Model. As illustrated in Fig. 5, we also compare our proposed SVGFusion with existing state-of-the-art approaches that directly generate SVGs using Large Language Models (LLMs). The results indicate that the performance of Claude3.5-sonnet [2], GPT4o-latest [1], and DeepSeek-R1 [11] in SVG generation is not particularly outstanding. In most cases, they can only use simple shapes to roughly assemble objects, but the positioning of each element lacks harmony. As a result, the overall shapes are overly simplistic, and the visual effects are less satisfactory. Regarding color design, these LLMs struggle to apply colors accurately to each part of the SVG, leading to color schemes that are often neither harmonious nor reasonable. In terms of semantic expression, the SVG code generated by LLMs is too simple to fully capture the meaning conveyed by the input text description. In contrast, our proposed SVGFusion method produces more balanced and harmonious results in terms of shape selection, color matching, and semantic representation. In Supplementary Sec. E, we provide more comparison of our method with language model-based methods.
@@
Editbility of SVGFusion Results. Figure 7 shows the editability of the SVG generated by our SVGFusion. Since the SVGs we generate have a clean and concise structure, we can easily edit the properties of the primitives, such as their color attributes. For instance, the rocket we generated can be changed from blue to red, or even transformed into a cyber-metal style, simply by adjusting the color attributes. Furthermore, the capability of edibility empowers users to efficiently reuse synthesized vector elements and create new vector compositions. As illustrated in the second example of Fig. 7, our method composes a new vector graphic by replacing the torch with an umbrella.